One number is two more than a second number. Twice the first is 4 less than 3 times the second. Find the numbers.

Answer by stanbon(75887)

One number is two more than a second number. Twice the first is 4 less than 3 times the second. Find the numbers.
----
Equations:
x = y+2
2x= 3y-4
----------
Substitute for x and solve for "y":
2(y+2) = 3y-4
2y+4 = 3y-4
y = 8 (the 2nd number)
-----
Since x = y + 2, x = 10 (the 1st number)
==============================
Cheers,
Stan H.